please show me the work 2x-4-5x=-x+5+7x-3
Answer by london maths tutor(243)

2x-4-5x=-x+5+7x-3
(2-5)x-4 = (-1+7)x + 5 - 3
(-3)x - 4 = 6x + 2
6x - 3x = 2 + 4
3x = 6
x = 6/3
x = 2